catalina:

Sam wants to buy 8 pencils and 6 pens, which cost a total of $8.50. He realizes he doesn't have enough cash and instead takes only 4 of each, which costs only $5. What is the price of each pen and pencil?

Hero:

Let x represent the cost of a pencil and y represent the cost of a pen Then 8x + 6y = 8.50 4x + 4y = 5 The system can be reduced to 4x + 3y = 4.25 4x + 4y = 5